OpenStudy (diana.xl):

It takes Nate four hours to paint a wall, and it takes Jill seven hours to paint the same wall. About how many hours will it will take if Nate and Jill work together to paint the wall? 11 hours 5.5 hours 2.5 hours 1.5 hours

OpenStudy (astrophysics):

@ParthKohli

Parth (parthkohli):

in an hour, Nate does 1/4th of the work and Jill does 1/7th of the work. If both work together, they'll do (1/4+1/7)th of the work in an hour. 1/4 + 1/7 = 11/28th of the work. So they'll require 28/11 hours.

OpenStudy (astrophysics):

ty

Parth (parthkohli):

dogs4dogs

OpenStudy (diana.xl):

so would my answer be 2.5 hours?

OpenStudy (welshfella):

here you can use the rate/hour of painting the wall Nate paints 1/4 of the wall in one hour, Jill paints 1/7 :- 1/4 + 1/7 = 1/x where x is the time if they work together

OpenStudy (anonymous):

crap you guys both beat me

OpenStudy (anonymous):

i was busy helping someone else @Astrophysics @ParthKohli

OpenStudy (welshfella):

the answer is 28/11 = 2 6/11 hours or 2.55 hours to nearest hundredth

Directrix (directrix):

The equation of @welshfella is correct.

Directrix (directrix):

>28/11 = 2 6/11 hours or 2.55 hours Or, 2 hours and 33 minutes, approximately @Diana.xL